# Audio-Language Models for Audio-Centric Tasks: A survey

### v1

@@ IV-C3 Discriminative Pre-trained Models
Discriminative pre-training aims to set up a pretext task of audio-text matching, allowing the model to learn cross-modal alignment features. MINT [34] is a framework that enhances audio-language pre-training through multiple objectives. Specifically, it introduces Bridge-Net as a trainable module, taking the output of the audio encoder and text as inputs to the Bridge-Net. It uses contrastive objective to align the audio and text representations by maximizing the mutual information between them, combines matching objective for fine-grained audio-text alignment, and employs generative objective to guide the audio-grounded text generation task, forcing the model to extract audio features to capture all necessary information about generating the text.
@@ V Downstream Transfer
Downstream transfer is crucial for enhancing the performance of ALMs, enabling them to adapt efficiently to new tasks. Transfer focuses on applying the knowledge gained from solving pretext tasks to solve other related tasks [1]. This can be divided into task-specific fine-tuning or multi-task tuning and cooperation.
@@ V-A Task-specific Fine-tuning
Common discriminative tasks include audio classification (AC) and audio-text retrieval (ATR). Audio classification broadly encompasses more specific scenarios such as Sound Event Classification, Accent Classification, Chord Classification, Spoken Language Identification, speech emotion recognition (SER), and spoken language understanding (SLU). Generative tasks can be further divided into audio generation and language generation. Typical audio generative tasks include: speech continuation (SC), text-to-speech (TTS), and speech enhancement (SE) in the speech domain [32, 84, 85, 86, 87, 88, 89]; piano continuation (PC), stereo generation (SG), and text-to-music (TTM) in the music domain [90, 91, 89]; as well as text-guided audio continuation (TAC), text-to-audio (TTA), language-query sound source separation (LASS), and text-query sound event detection (TSED) in other general audio domains [92, 93, 4, 6, 94, 95, 96, 97, 98]. Language generation, on the other hand, centers on transcription, translation, and caption. Typical tasks include automatic speech recognition (ASR) [47, 99, 100], speech-to-text translation (S2TT) [84], speech-to-speech translation (S2ST) [47], and automatic audio captioning (AAC) [101, 102, 5, 103, 104, 105].
@@
Model Domain Architectures / Pre-trained models Tasks Input Output
@@ A T A T
VALL-E [32] Speech EnCodec+AR/NAR LM TTS
VALL-E X [84] EnCodec+AR/NAR LM TTS, S2ST
MMCE-Qformer [85] RoBERTa+SpeechTokenizer+GPT-2 TTS
E2 TTS [86] Unsupervised Pretraining+BigVGAN TTS
Seed-TTS [87] Speech Tokenizers+AR LM+DiT+Vocoders TTS
F5-TTS [88] ConvNeXt+DiT TTS
Whisper [47] BPE + encoder-decoder Transformer ASR, S2TT
Style-Talker [99] Whisper+Qwen-Audio+StyleTTS 2 ASR
Seed-ASR [100] LUISE+MoE LLM ASR
MusicLM [90] Music SoundStream+w2v-BERT+MuLan TTM
MusicGen [106] EnCodec+AR LMs TTM, SG
MeLoDy [107] MuLan+Wav2Vec2-Conformer+VAE+LM+DPD TTM
TANGO [91] FLAN-T5+LDM+VAE+HiFi-GAN TTM
LP-MusicCaps [101] Whisper+BART-base AAC
AudioLM [89] Speech,Music SoundStream+w2v-BERT SC, PC
AudioGEN [92] Audio SoundStream+LSTM TAC
Diffsound [93] VQ-VAE+diffusion TTA
AudioLDM [4] CLAP+LDM+VAE TTA
LASS-Net [6] BERT+ResUNet LASS
ClipSep [94] CLIP+U-Net LASS
AudioSep [95] CLIP/CLAP+ResUNet LASS, SE
OpenSep [96] Caption Models+Instuction tuned LLMs LASS
FlowSep [97] FLAN-T5+VAE+BigVGAN LASS
TQ-SED [98] CLAP+ResUNet+DPRNN TSED
ACT [102] DeiT+Word2Vec AAC
EnCLAP [5] CLAP+BART+EnCodec AAC
EnCLAP++ [103] CLAP+BART+DAC AAC
ReCap [104] CLAP+GPT-2 AAC
DRCap [105] CLAP+SLAM LLM AAC
  • Please refer to the main text for abbreviations.
@@ V-A1 Models for Discriminative Tasks
Through contrastive learning with natural language supervision, audio-language pre-trained models can perform zero-shot audio classification and retrieval tasks within a unified framework, as shown in Fig. 6. Although pre-trained models already have strong generalization capabilities, supervised fine-tuning on downstream datasets remains an important step to enhance task performance [2, 22, 50]. Audio classification refers to mapping input audio to a label, and the audio encoder from the pre-trained model can serve as a powerful audio pre-training model, used to train additional classifiers on specific datasets for fine-tuning. Retrieval can be divided into audio-to-text and text-to-audio, aiming to filter corresponding samples of another modality from a pool based on the input, with classification considered as a special case of retrieval.
ATR research primarily focuses on enhancing system performance by integrating orthogonal SOTA works. For instance, CNN14-NetRVLAD [108] explores various audio features and sequence aggregation methods to improve audio-text alignment, utilizing the CNN14-based audio pretraining model PANNs and NetRVLAD pooling techniques to achieve significant improvements in bidirectional audio-text retrieval. EN-CA-IMC [109] implements an audio enhancement strategy (EN) involving Gaussian noise, pitch adjustment, and time shift to reduce model overflow, combined with a co-attentive mechanism (CA) and intra-modal contrastive learning (IMC) between enhanced and original audio to capture richer audio features. CED-BERT-LE [110] integrates a language enhancement strategy (LE) for effective cross-lingual retrieval and uses consistent ensemble distillation (CED) to handle variable-length audio segments, demonstrating proficiency across seven additional languages with minimal extra training data. Recently, the mini-batch Learning-to-match (m-LTM) framework [111] adapts the Learning-to-match approach using mini-batch subsampling and Mahalanobis-enhanced ground metrics, leveraging soft-matching from entropic optimal plans and Mahalanobis distance to learn a rich, expressive joint embedding space for audio and text modalities.
@@ V-A2 Models for Generative Tasks
Audio generation aims to enable the model to learn to generate audio that better conforms to textual conditions, facilitating the creation of personalized audio for applications such as augmented reality and virtual reality, game development, and video editing [4]. A common method is to use a codec structure that converts continuous audio into discrete tokens for language modeling, ensuring high fidelity [52]. AudioGEN [92] is an pioneering audio language codec that uses data augmentation techniques, multi-stream modeling, and classifier-free guidance to improve adherence to text, addressing challenges such as distinguishing multiple sound sources, handling real-world recording conditions, and encoding high-fidelity audio. VALL-E [32], VALL-E X [84], and MusicLM [90] perform well in speech and music respectively. SpeechX [112] combines neural codec language modeling with task-specific prompts to achieve unified and scalable modeling and provides a consistent method for utilizing text input in speech enhancement and conversion tasks. Recent approaches further advance the development of audio generation technology by combining advanced generative models such as flow models and Diffusion. flow-based models like Flow-TTS [113] improve the overall fluency and naturalness of generated speech by using reversible transformations to balance local accuracy and global coherence. Diffsound [93] is an work based on diffusion models, using text as a condition for controlled audio generation by the diffusion model. Seed-TTS [87] is a foundational speech generation model that excels in speech in-context learning and achieves good performance in speaker similarity and naturalness.
Audio source separation is a crucial task in audio enhancement applications and can be considered a type of audio generative task according to objectives in fine-tuning. LASS-Net [6] is the first proposed end-to-end network for the language-queried audio source separation (LASS) task, capable of separating the target source from an audio mixture based on user’s natural language descriptions (e.g., “a woman speaking”). Transfer ALMs to such tasks typically requires new design in the model architecture. For example, as shown in Fig. 7, AudioSep [95] uses the text pretraining that achieves audio-text modality alignment during the ALM pretraining process as the query encoder for text and connects it to the prediction network for downstream tasks (audio separation network), and conduct post-training on specific task datasets. It not only achieved audio segmentation for arbitrary descriptions but was also extended to speech enhancement tasks, demonstrating the flexibility of combining natural language prompts with task paradigms. This provides the potential to integrate multiple audio enhancement tasks into a unified framework. Current research is unfolding in two main aspects: improving separation performance [97] and enhancing generalizability [96].
@@
Language generation aims to enable the model to perform audio-centric text prediction tasks, and fine-tuning can be conducted through language generative objectives. The two sub-tasks, audio understanding and language generation, are accomplished by the audio encoder and the text decoder, respectively. Recent work has also built upon this basic architecture by adding multiple encoders or decoders [17].
Automated Audio Captioning (AAC) is an emerging audio perception task that has gained popularity in recent years. It involves recognizing environments, sound events, and the temporal relationships between these events, and then describing them in fluent sentences. Currently, the standard approach to address this problem is based on an encoder-decoder deep learning framework, with many studies focusing on innovative network architectures and training schemes for improvements [16]. The Audio Captioning Transformer (ACT) [102] is the first full Transformer architecture designed specifically for AAC, excelling in capturing global information and temporal relationships within audio signals. EnCLAP [5] integrates two powerful acoustic representation models, EnCodec and CLAP, with the pre-trained language model BART to generate more accurate and contextually relevant captions. AutoCap [114] leverages extra guidance from metadata to enhance feature extraction, capturing nuanced acoustic information. LP-MusicCaps [101] addresses data scarcity in music captioning by using LLMs to generate descriptions from large-scale tag datasets.
@@ V-B Multi-task Tuning and Cooperation
In the early stages of audio processing, research efforts were primarily focused on breaking down complex tasks into simpler, independent audio tasks. Some studies [1, 115] suggest that while these task-specific models were effective, they lacked understanding capabilities and failed to capture the connections between tasks, resulting in poor scalability and limited robustness when dealing with the inherent variability and complexity of audio signals. Some remarkable work has combined LLM in ALM achieving understanding and reasoning across diverse audio tasks. This convergence of large-scale language modeling with audio processing has opened up new possibilities to create more powerful, context-aware, and human-like audio understanding and reasoning systems.
@@
Type Model Domain Architectures / Pre-trained models Input Output
A T A T
Codec UniAudio [116] Audio Tokenizers+Multi-scale Transformer
SpeechX [112] Speech EnCodec+AR/NAR LM (VALL-E)
VioLA [117] LSTM+AR LM
AudioPaLM [7] w2v-BERT/USM-v1/USM-v2+PaLM-2+SoundStream
LauraGPT [118] EnCodec+Qwen+Vododer
LALM Pengi [1] Audio CLAP+GPT-2
LTU [24] AST+LLaMA
SALMONN [119] Q-Former+Whisper+BEATs+Vicuna
Qwen-audio [66] Whisper+Qwen
Audio Flamingo [120] CLAP+OPT-IML-MAX
GAMA [121] AST+Q-Former+LLaMa-2
SpeechGPT [8] Speech HuBERT+LLaMA+HiFi-GAN
SALM [122] Conformer+Megatron LLM
BLSP [123] Whisper+Llama-2
SpeechT5 [124] wav2vec 2.0+Transformer+vocoders
Voicebox [125] ALiBi+VB-En/VB-Multi
SpeechVerse [115] WavLM/Best-RQ/Others+Flan-T5
COSMIC [126] Whisper-medium+Qformer+LLaMA-2
Moshi [127] Mimi+Helium
MusiLingo [128] Music MERT+Vicuna
Llark [129] Jukebox+Llama-2
Agent System AudioGPT [130] Audio GPT-3.5+Audio Foundation Models
SpeechAgents [131] Speech SpeechGPT+LLaMA-2
MusicAgents [132] Music LLMs+ Music Tools
As listed in Table III, the recently popular ALMs can be categorized into two main types: Multi-task Models (including codec-based models and LALMs) and Audio-Language Agent Systems.
@@ V-B1 Multi-task Tuning Models
Multi-task sequence-to-sequence modeling is a classic approach involves designing a multi-task sequence format for jointly represented tokens, which are widely adopted by codec-based ALMs. Codecs convert continuous audio into discrete tokens, which can easily be merged with text, MIDI, and other contextual tokens, and then passed to a language model to generate an output sequence that is further converted into the desired audio, text, or other outputs as needed for the task [14]. This codec-based pipeline based on discrete token has the advantage of high fidelity and is widely used in generative tasks. Moreover, its flexible encoding allows for more natural interaction between text and audio, enabling the use of a single model to perform ASR, TTS, S2TT, MT, and S2ST tasks, using context or ID to indicate the task to which a sample belongs [117]. Depending on the type of language model employed, codecs can be categorized into autoregressive (AR) and non-autoregressive (NAR) models. However, they may suffer from information loss due to the quantization of voice signals into discrete tokens, leading to significant performance degradation in some tasks compared to models that use continuous speech features [118]. Apart from codecs, Large Audio-Language Models (LALMs) aims to leverage the emergent understanding and reasoning capabilities of LLMs to build more generalized audio-centric models. These models typically employ an audio pre-trained model as the encoder, set up encoders for other modalities as needed, and add adaptive modules before connecting an LLM on top. However, pre-trained LLMs still lack sufficient cross-task generalization capabilities and require further post-training [11]. Multi-task tuning for LALMs can be similar to codecs, adopting a unified sequence format, or achieved through instruction tuning or in-context learning. VioLA [117] is a Transformer-based autoregressive decoder network that unifies cross-modal tasks involving speech and text. The model converts speech into discrete tokens and integrates task IDs and language IDs to enhance the ability to handle different languages and tasks. Experimental results demonstrate its excellent performance in both single-modal and cross-modal tasks. LauraGPT [118] operates based on a unified task expression (input embeddings, task ID, output tokens) to handle multiple tasks. Its modular and flexible design allows the model to perform complex tasks by combining sub-tasks such as S2ST. AudioPaLM [7] is a model for speech understanding and generation. In combined task learning, the model is instructed to output intermediate steps of complex tasks, which aligns with the spirit of chain-of-thought prompting. For example, in Speech-to-Speech Translation (S2ST), the model can perform end-to-end conversion from English audio tokens to French audio tokens through a single autoregressive decoding process, guided by task tags like [ASR AST S2ST English French]. This enables the integration of ASR, MT, and TTS within a unified model, eliminating the need for a separated pipeline approach. UniAudio [116] is an advanced codec-based ALM supports 11 diverse audio generation tasks. It achieves this by tokenizing both target audio and various condition modalities, then concatenating these source-target pairs into unified [conditions, target] sequences. The system leverages LM for next-token prediction and incorporates a multi-scale Transformer architecture to efficiently handle the extended sequence lengths introduced by neural codec-based tokenization. Pengi [1] frames all audio tasks as text-generation tasks, allowing for comprehensive audio understanding and generation. Pengi demonstrates remarkable capabilities across zero-shot tasks in general audio including environmental sound, speech, and music tasks, and achieves good performance in close-ended and open-ended audio tasks. Qwen-audio [33] adopts a multi-task training framework, which uses hierarchical tags to condition the decoder, allowing for effective knowledge sharing while avoiding interference issues. This approach enables the model to handle diverse audio types and tasks, including automatic speech recognition (ASR), audio captioning, and audio question-answering. Qwen-audio is pre-trained across over 30 tasks in environmental sound, speech, and music, and achieves state-of-the-art results on 9 evaluation benchmarks without requiring task-specific fine-tuning, surpassing baselines on several tasks such as automatic speech recognition and acoustic scene classification.

### v2

@@ IV-C3 Discriminative Pre-trained Models
Discriminative pre-training aims to set up a pretext task of audio-text matching, allowing the model to learn cross-modal alignment features. MINT [zhao2024mint] is a framework that enhances audio-language pre-training through multiple objectives. Specifically, it introduces Bridge-Net as a trainable module, taking the output of the audio encoder and text as inputs to the Bridge-Net. It uses contrastive objective to align the audio and text representations by maximizing the mutual information between them, combines matching objective for fine-grained audio-text alignment, and employs generative objective to guide the audio-grounded text generation task, forcing the model to extract audio features to capture all necessary information about generating the text.
@@ V Downstream Transfer
Downstream transfer is crucial for enhancing ALM performance and enabling efficient adaptation to new tasks. It involves applying knowledge acquired from pretext tasks to solve related problems [deshmukh2023pengi]. This process adapts general-purpose models to domain-specific needs, facilitating deployment of robust systems in real-world applications like voice assistants and audio search. Transfer methods comprise task-specific fine-tuning and multi-task tuning/cooperation.
@@ V-A Task-specific Fine-tuning
Common discriminative tasks include audio classification (AC) and audio-text retrieval (ATR). Audio classification encompasses specific scenarios such as sound event detection, accent recognition, chord identification, spoken language identification, speech emotion recognition (SER), and spoken language understanding (SLU). Generative tasks are categorized into audio generation and language generation. Typical audio generative tasks include: speech continuation (SC), text-to-speech (TTS), and speech enhancement (SE) in the speech domain [wang2023VALLE, zhang2023VALLEX, xue2024mmace, eskimez2024e2tts, anastassiou2024seedtts, chen2024f5tts, Borsos2023audiolm]; piano continuation (PC), stereo generation (SG), text-to-music (TTM) and music editing (ME) in the music domain [agostinelli2023musiclm, melechovsky2024mustango, Borsos2023audiolm]; as well as text-guided audio continuation (TAC), text-to-audio (TTA), text-to-spatial-audio (TTSA, also referred as spatial audio generation), language-query sound source separation (LASS), and text-to-audio grounding (TAG, also referred as text-query sound event detection) in other general audio domains [kreuk2023audiogen, yang2023diffsound, liu2023audioldm, liu2022_lass, dong2023clipsep, liu2024separate, mahmud2024opensep, yuan2025flowsep, yin2024TQSED, Xu2024TAG]. Language generation, centers on transcription, translation, and caption. Typical tasks include automatic speech recognition (ASR) [radford2023whisper, li2024style, bai2024seedasr], speech-to-text translation (S2TT) [zhang2023VALLEX], speech-to-speech translation (S2ST) [radford2023whisper], and automatic audio captioning (AAC) [doh2023lp, mei2021audio, kim2024enclap, Kim2024enclap++, Ghosh2024Recap, li2024drcap].
@@
[…]
@@ A T A T
[…]
VALL-E X [zhang2023VALLEX] EnCodec+AR/NAR LM TTS, S2ST
E2 TTS [eskimez2024e2tts] Unsupervised Pretraining+BigVGAN TTS
Seed-TTS [anastassiou2024seedtts] Speech Tokenizers+AR LM+DiT+Vocoders TTS
F5-TTS [chen2024f5tts] ConvNeXt+DiT TTS
Whisper [radford2023whisper] BPE + encoder-decoder Transformer ASR, S2TT
Style-Talker [li2024style] Whisper+Qwen-Audio+StyleTTS 2 ASR
Seed-ASR [bai2024seedasr] LUISE+MoE LLM ASR
LLaSE-G1 [kang2025llaseg1] WavLM+LLaMA+Xcodec2 SE
MusicLM [agostinelli2023musiclm] Music SoundStream+w2v-BERT+MuLan TTM
MusicGen [copet2024simple] EnCodec+AR LMs TTM, SG
TANGO [melechovsky2024mustango] FLAN-T5+LDM+VAE+HiFi-GAN TTM
LP-MusicCaps [doh2023lp] Whisper+BART-base AAC
SongEditor [yang2025songeditor] RVQ+LM+DiT TTM, ME
AudioLM [Borsos2023audiolm] Speech, Music SoundStream+w2v-BERT SC, PC
AudioGEN [kreuk2023audiogen] Audio SoundStream+LSTM TAC
Diffsound [yang2023diffsound] VQ-VAE+diffusion TTA
AudioLDM [liu2023audioldm] CLAP+LDM+VAE TTA
DualSpec [zhao2025dualspec] VAE+Flan-T5+LDM TTSA
ClipSep [dong2023clipsep] CLIP+U-Net LASS
AudioSep [liu2024separate] CLIP/CLAP+ResUNet LASS, SE
OpenSep [mahmud2024opensep] Caption Models+Instuction tuned LLMs LASS
FlowSep [yuan2025flowsep] FLAN-T5+VAE+BigVGAN LASS
TQ-SED [yin2024TQSED] CLAP+ResUNet+DPRNN TSED
WSTAG [Xu2024TAG] CRNN+CLAP TAG
ACT [mei2021audio] DeiT+Word2Vec AAC
EnCLAP [kim2024enclap] CLAP+BART+EnCodec AAC
EnCLAP++ [Kim2024enclap++] CLAP+BART+DAC AAC
DRCap [li2024drcap] CLAP+SLAM LLM AAC
[…]
@@ V-A1 Models for Discriminative Tasks
Through contrastive learning with natural language supervision, audio-language pre-trained models can perform zero-shot audio classification and retrieval tasks within a unified framework, as shown in Fig. 6. Although pre-trained models already have strong generalization capabilities, supervised fine-tuning on downstream datasets remains an important step to enhance task performance [elizalde2023clap, wu2023large, elizalde2024CLAP2]. Audio classification refers to mapping input audio to a label, and the audio encoder from the pre-trained model can serve as a powerful audio pre-training model, used to train additional classifiers on specific datasets for fine-tuning. Retrieval can be divided into audio-to-text (A-T) and text-to-audio (T-A), aiming to filter corresponding samples of another modality from a pool based on the input, with classification considered as a special case of retrieval.
ATR research primarily focuses on enhancing system performance by integrating orthogonal SOTA works. For instance, CNN14-NetRVLAD [lou2022contextatr] explores various audio features and sequence aggregation methods to improve audio-text alignment, utilizing the CNN14-based audio pretraining model PANNs and NetRVLAD pooling techniques to achieve significant improvements in bidirectional audio-text retrieval. EN-CA-IMC [hu2023enatr] implements an audio enhancement strategy (EN) involving Gaussian noise, pitch adjustment, and time shift to reduce model overflow, combined with a co-attentive mechanism (CA) and intra-modal contrastive learning (IMC) between enhanced and original audio to capture richer audio features. CED-BERT-LE [yan2024bridging] integrates a language enhancement strategy (LE) for effective cross-lingual retrieval and uses consistent ensemble distillation (CED) to handle variable-length audio segments, demonstrating proficiency across seven additional languages with minimal extra training data. Recently, the mini-batch Learning-to-match (m-LTM) framework [luong2024revisiting] adapts the Learning-to-match approach using mini-batch subsampling and Mahalanobis-enhanced ground metrics, leveraging soft-matching from entropic optimal plans and Mahalanobis distance to learn a rich, expressive joint embedding space for audio and text modalities.
@@ V-A2 Models for Generative Tasks
Audio generation aims to enable the model to learn to generate audio that better conforms to textual conditions, facilitating the creation of personalized audio for applications such as augmented reality and virtual reality, game development, and video editing [liu2023audioldm].
@@
A common method is to use a codec structure that converts continuous audio into discrete tokens for language modeling, ensuring high fidelity [wu2024towards]. AudioGEN [kreuk2023audiogen] is an pioneering audio language codec that uses data augmentation techniques, multi-stream modeling, and classifier-free guidance to improve adherence to text, addressing challenges such as distinguishing multiple sound sources, handling real-world recording conditions, and encoding high-fidelity audio. VALL-E [wang2023VALLE], VALL-E X [zhang2023VALLEX], MusicLM [agostinelli2023musiclm] and SongEditor [yang2025songeditor] perform well in speech and music respectively. SpeechX [wang2024speechx] combines neural codec language modeling with task-specific prompts to achieve unified and scalable modeling and provides a consistent method for utilizing text input in speech enhancement and conversion tasks. Recent approaches further advance the development of audio generation technology by combining advanced generative models such as flow models and Diffusion. Flow-based models like Flow-TTS [miao2020flow] improve the overall fluency and naturalness of generated speech by using reversible transformations to balance local accuracy and global coherence. Extending this approach, UniFlow-Audio [xu2025uniflow] establishes a universal non-autoregressive framework for general audio generation, handling both time-aligned and non-time-aligned tasks via a dual-fusion mechanism that integrates features through temporal alignment and cross-attention. Diffsound [yang2023diffsound] is an work based on diffusion models, using text as a condition for controlled audio generation by the diffusion model. Seed-TTS [anastassiou2024seedtts] is a foundational speech generation model that excels in speech in-context learning and achieves good performance in speaker similarity and naturalness.
Audio source separation is a crucial audio generation task in enhancement applications. The language-queried audio source separation (LASS) task, which aims to isolate a target sound from a mixture based on natural language descriptions (e.g., “a woman speaking”), typically requires tailored architectures when transferring ALMs. A representative approach is AudioSep [liu2024separate], which leverages a pre-trained ALM with aligned audio-text representations as its text query encoder. As illustrated in Fig. 7, this encoder connects to a separation network and is fine-tuned on task-specific data, enabling source separation from free-form text queries and extending to tasks like speech enhancement. This demonstrates the potential of using natural language prompts to unify various audio enhancement tasks. Current research focuses on improving separation quality [yuan2025flowsep] and model generalizability [mahmud2024opensep].
Recent works have explored spatial audio generation, advancing the field from single-channel to multi-channel and from non-directional to controllable formats. DualSpec [zhao2025dualspec] is the first framework to directly generate spatial stereo audio with directional information from text descriptions, through a dual-branch design that integrates Mel-spectrograms and short-time Fourier transform spectrograms. SALM [hu2025salm] aligns spatial audio with language via multi-modal contrastive learning. Its dual-branch encoder decomposes audio into semantic and spatial components, producing structured embeddings that enable zero-shot direction classification and text-based spatial audio editing. These technological advancements provide prospects for constructing more realistic audio datasets through synthetic data augmentation methods.
@@
Language generation aims to enable the model to perform audio-centric text prediction tasks, and fine-tuning can be conducted through language generative objectives. The two sub-tasks, audio understanding and language generation, are accomplished by the audio encoder and the text decoder, respectively. Recent work has also built upon this basic architecture by adding multiple encoders or decoders [xu2023S2TT].
Automated Audio Captioning (AAC) is an emerging task that involves recognizing environments, sound events, and their temporal relationships, then describes them in fluent sentences. The standard framework uses encoder-decoder architectures, with recent studies focusing on novel networks and training schemes for improvements[xu2023aac_survey]. The Audio Captioning Transformer (ACT) [mei2021audio] is the first full Transformer architecture designed specifically for AAC, excelling in capturing global information and temporal relationships within audio signals. EnCLAP [kim2024enclap] integrates two powerful acoustic representation models, EnCodec and CLAP, with the pre-trained language model BART to generate more accurate and contextually relevant captions. AutoCap [haji2024taming] leverages extra guidance from metadata to enhance feature extraction, capturing nuanced acoustic information. LP-MusicCaps [doh2023lp] addresses data scarcity in music captioning by using LLMs to generate descriptions from large-scale tag datasets.
@@ V-B Multi-task Tuning and Cooperation
In the early stages of audio processing, research efforts were primarily focused on breaking down complex tasks into simpler, independent audio tasks. Some studies [deshmukh2023pengi, das2024speechverse] suggest that while these task-specific models were effective, they lacked understanding capabilities and failed to capture the connections between tasks, resulting in poor scalability and limited robustness when dealing with the inherent variability and complexity of audio signals. Some remarkable work has combined LLM in ALM achieving understanding and reasoning across diverse audio tasks. This convergence of large-scale language modeling with audio processing has opened up new possibilities to create more powerful, context-aware, and human-like audio understanding and reasoning systems.
@@
[…]
Codec UniAudio [yang2024uniaudio] Audio Tokenizers+Multi-scale Transformer
SpeechX [wang2024speechx] Speech EnCodec+AR/NAR LM (VALL-E)
[…]
LALM Pengi [deshmukh2023pengi] Audio CLAP+GPT-2
[…]
SpeechGPT [zhang2023speechgpt] Speech HuBERT+LLaMA+HiFi-GAN
[…]
MusiLingo [deng2024musilingo] Music MERT+Vicuna
[…]
Agent System AudioGPT [huang2024audiogpt] Audio GPT-3.5+Audio Foundation Models
SpeechAgents [zhang2024speechagents] Speech SpeechGPT+LLaMA-2
MusicAgents [yu2023musicagent] Music LLMs+ Music Tools
[…]
As listed in Table III, the popular ALMs can be categorized into two types: Multi-task Models (including codec-based models and LALMs) and Audio-Language Agent Systems.
@@ V-B1 Multi-task Tuning Models
Multi-task sequence-to-sequence modeling is a classic approach involves designing a multi-task sequence format for jointly represented tokens, which are widely adopted by codec-based ALMs. Codecs convert continuous audio into discrete tokens, which can easily be merged with text, MIDI, and other contextual tokens, and then passed to a language model to generate an output sequence that is further converted into the desired audio, text, or other outputs as needed for the task [wang2023codec]. This codec-based pipeline based on discrete token has the advantage of high fidelity and is widely used in generative tasks. Moreover, its flexible encoding allows for more natural interaction between text and audio, enabling the use of a single model to perform ASR, TTS, S2TT, MT, and S2ST tasks, using context or ID to indicate the task to which a sample belongs [wang2023viola]. Depending on the type of language model employed, codecs can be categorized into autoregressive (AR) and non-autoregressive (NAR) models. However, they may suffer from information loss due to the quantization of voice signals into discrete tokens, leading to significant performance degradation in some tasks compared to models that use continuous speech features [du2023lauragpt]. Apart from codecs, Large Audio-Language Models (LALMs) aims to leverage the emergent understanding and reasoning capabilities of LLMs to build more generalized audio-centric models. These models typically employ an audio pre-trained model as the encoder, set up encoders for other modalities as needed, and add adaptive modules before connecting an LLM on top. However, pre-trained LLMs still lack sufficient cross-task generalization capabilities and require further post-training [zhao2023LLMsurvey]. Multi-task tuning for LALMs can be similar to codecs, adopting a unified sequence format, or achieved through instruction tuning or in-context learning. VioLA [wang2023viola] is a Transformer-based autoregressive decoder that unifies speech-text cross-modal tasks. By converting speech to discrete tokens and integrating task/language IDs, it effectively handles multiple languages and tasks, demonstrating excellent performance in both single-modal and cross-modal settings. LauraGPT [du2023lauragpt] operates based on a unified task expression (input embeddings, task ID, output tokens) to handle multiple tasks. Its modular and flexible design allows the model to perform complex tasks by combining sub-tasks such as S2ST. UniAudio [yang2024uniaudio] is an advanced codec-based ALM that supports 11 diverse audio generation tasks. It tokenizes both target audio and various condition modalities, concatenating them into unified [conditions, target] sequences. The system employs a multi-scale Transformer architecture for next-token prediction, efficiently handling the extended sequences from neural codec tokenization. Pengi [deshmukh2023pengi] frames all audio tasks as text-generation tasks, allowing for comprehensive audio understanding and generation. Pengi demonstrates remarkable capabilities across zero-shot tasks in general audio including environmental sound, speech, and music tasks, and achieves good performance in close-ended and open-ended audio tasks. Qwen-audio [chu2023qwenaudio] employs a multi-task training framework that uses hierarchical tags to condition the decoder, enabling effective knowledge sharing without interference. This allows the model to handle diverse audio types and tasks, including ASR, AAC, and audio question answering. SHANKS [chiang2025shanks] and STITCH [chiang2025stitch] exemplify this through specialized training sequences that interleave input, reasoning, and output. SHANKS interleaves user speech chunks with internal ”thinking” tokens to learn incremental reasoning for low-latency intervention. STITCH alternates between reasoning and speech output tokens, enabling concurrent ”thinking” and ”speaking”.